OpenStudy (anonymous):

A rectangle has an area of 24 square centimeters and a length of 3 centimeters. What is the width of the rectangle? A.4 centimeters B.6 centimeters C.8 centimeters D.21 centimeters

OpenStudy (anonymous):

A = L*W 24 = 3*W divide both sides by 3 w = 8

OpenStudy (anonymous):

thx